A boy and a girl work together to type a thesis and it takes them 6 hours. It would take the girl 5 hours more than the boy to type the thesis alone. How long would each need to type the thesis alone?
Answer by fractalier(6550)

Let the time it takes the boy alone to be x. Her time is then x + 5.
The setup for problems like this is
t/A + t/B = 1
where t is the time worked together and A and B are the individual times...so plugging in we get
6/x + 6/(x+5) = 1
now multiply thru by the lowest common denominator, here x(x+5)...and we get
x(x+5)[6/x + 6/(x+5) = 1]
6(x+5) + 6x = x(x+5)
6x + 30 + 6x = x^2 + 5x
now rearrange, factor and solve
x^2 - 7x - 30 = 0
(x - 10)(x + 3) = 0
x = 10 or x = -3
since it cannot be negative time,
x = 10 hours and
x+5 = 15 hours